# Evaluating the Effect of Retrieval Augmentation on Social Biases

### v1

@@ Abstract
Llm-jp-13B 17.96 / 6.55 23.02 / 15.67 3.08 / 2.58 6.35 / -0.79
Qwen-3B 28.27 / 8.13 39.83 / 8.13 24.70 / -1.59 11.81 / -6.15
Qwen-3B-Inst. 17.41 / 0.20 23.86 / 4.07 15.18 / -5.75 6.35 / -8.93
Qwen-7B 18.85 / -1.39 27.88 / 0.00 17.91 / -3.97 10.02 / -8.63
Qwen-7B-Inst. 10.02 / -3.67 24.01 / 0.50 15.43 / -2.08 10.17 / -10.12
Qwen-14B 3.77 / -7.34 13.99 / -2.68 0.55 / -4.66 -4.51 / -8.93
Qwen-14B-Inst. -2.38 / -2.38 4.61 / 2.68 -3.08 / -5.95 -8.43 / -12.70
@@
Table 3: Diff-Bias scores for the ambiguous and disambiguated gender contexts (separated by ‘/’) for different generator LLMs. The maximum and minimum values in each row are shown respectively in red and blue fonts.
Model CBBQ JBBQ
w/o RAG stereo-set full-set anti-set w/o RAG stereo-set full-set anti-set
GPT-3.5 18.07 / 8.64 35.61 / 16.26 13.74 / 6.79 -6.39 / 6.38 1.51 / -4.75 12.17 / 2.15 11.50 / 1.84 3.25 / 0.31
Qwen-7B-Inst. 7.79 / 3.91 46.00 / 23.05 25.43 / 12.76 -4.33 / -6.58 1.53 / -5.06 13.11 / -7.21 10.35 / -5.98 10.53 / -4.65
Qwen-14B 9.85 / -0.62 32.47 / 13.17 7.25 / 0.00 -6.60 / -10.70 8.77 / -16.00 17.41 / -9.36 11.58 / -12.93 9.56 / -17.94
Qwen-14B-Inst. 3.68 / 1.44 21.97 / 17.49 6.39 / -4.12 -9.09 / -13.58 -0.72 / -20.50 11.84 / -19.22 4.22 / -15.59 3.73 /-19.79
Table 4: Diff-Bias scores for Chinese (CBBQ) and Japanese (JBBQ) datasets, reported in the format ambiguous / disambiguated. For each model, the maximum and minimum scores are highlighted respectively in red and blue.
§ 4.3 Effect of the Generators
@@
To assess how RAG impacts different generator LLMs, we measure their gender-related social biases in the English BBQ dataset (see Table 3). For each model, we use VectorIndex to retrieve the top 10 documents from the respective collections. Table 3 shows LLMs trained on multilingual pre-train data in the top block, while models that are trained on increased proportions of Japanese and Chinese language pre-train data are shown respectively in the middle and bottom blocks.
Overall, every model exhibits increased gender bias when retrieving from the full-set or stereo-set, and decreased bias when retrieving from the anti-set. These findings corroborate the trend noted in Table 2, highlighting how RAG can amplify social biases in both advantaged and disadvantaged groups. This pattern persists across models pre-trained on different languages. Furthermore, within the Qwen family, larger instruction-tuned models generally show lower levels of gender bias.
§ 4.4 Multilingual Bias Evaluation
To examine how RAG influences social biases in languages beyond English, we extend our experiments to Japanese and Chinese using the JBBQ and CBBQ datasets, respectively. Both datasets follow the same QA format and address identical social bias types as the English BBQ dataset, making them suitable benchmarks for our evaluations. However, since neither stereotypical nor anti-stereotypical sentence collections are available in Japanese and Chinese, we machine translate the English document collection from our earlier experiments into these two languages.
Table 4 presents the Diff-Bias scores on the CBBQ and JBBQ datasets. In Chinese, retrieving documents from the stereo-set consistently amplifies social biases relative to the w/o RAG baseline for the advantaged group, often to a greater degree than in English. One possible explanation is that the machine translation process may have introduced additional biases into the document collection. In contrast, the anti-set increases bias toward disadvantaged groups compared to w/o RAG for all LLMs. Interestingly, for GPT-3.5 and in ambiguous contexts for Qwen-14B, the full-set yields lower social bias than w/o RAG, possibly due to balancing effects from the anti-set documents.
For Japanese, we similarly observe a consistent rise in social bias when retrieving from the stereo-set, compared to the w/o RAG baseline. In the ambiguous contexts, stereo-set typically produces the largest bias in favour of advantaged groups. However, the anti-set has a less predictable impact than in English and Chinese. For instance, Qwen7B-Inst. exhibits even higher bias with anti-set than with stereo-set. A closer examination indicates that machine translation may fail to preserve certain nuances of the original stereotypes, and Japanese-specific issues such as zero-pronoun resolution Isozaki and Hirao (2003) (i.e. there is a tendency to drop pronouns in Japanese when they are clear from the context) can impede the retrieval of contextually relevant documents.
§ 4.5 Effect of the Retrievers
@@
Table 5: Comparison of ambiguous and disambiguated Diff-Bias scores (separated by ‘/’) when using different retrieval methods to retrieving documents from the full-set. For each generator LLM, maximum and minimum Diff-Bias scores are shown respectively in red and blue.
To assess how different retrieval methods affect social biases in RAG we experiment with three approaches: VectorIndex, BM25 and Contriever—using the English BBQ dataset. Specifically, we measure the gender-related Diff-Bias of various generator LLMs when retrieving 10 documents from the full-set in Table 5. The percentages of stereotypical documents among the documents retrieved by each method are shown in the first row (stereo docs). We see that VectorIndex retrieves more balanced number of documents (i.e. approximately 50% stereotypical) compared to Contriever and BM25. Despite this behaviour, we see that all retrieval methods tend to amplify Diff-Bias scores compared to w/o RAG. Although BM25 retrieves least percentage of stereotypical documents compared to Contriever and VectorIndex, it shows a high level of biases across LLMs. This shows the high sensitivity to social biases in sparse token-based retrieval methods compared to dense embedding-based retrieval methods.
We next explore how varying the number of retrieved documents influences bias by using VectorIndex for three generator LLMs as shown in Figure 3 and Figure 4 respectively for the ambiguous and disambiguated contexts. In both full-set and stereo-set, ambiguous Diff-Bias scores rise sharply even with a small number of retrieved documents, compared to w/o RAG. However. after retrieving five or more documents, Diff-Bias scores begin to decrease—particularly for the larger Qwen-14B model. A similar trend occurs in disambiguated contexts, as the absolute Diff-Bias values lessen with more documents retrieved, except in the anti-set scenario. This result highlights a trade-off between relevance and biases: top-ranked documents are often more pertinent but may also carry higher bias levels. Notably, the larger Qwen-14B model appears more capable of mitigating bias when provided with a larger pool of documents. Accuracy-based evaluations for all experiments are shown in Appendix C and overall lead to similar conclusions as the ones made using Diff-Bias scores.
@@
We conducted a comprehensive study on how RAG influences social biases LLM outputs. Using the BBQ benchmark across multiple languages—English, Japanese, and Chinese—we demonstrated that introducing a retrieval component can significantly amplify social biases in generated text, even for models that appear relatively unbiased when used in isolation.
Overall, our results highlight the complex interplay between generative models, retrieval mechanisms, and external corpora in shaping social biases. We urge practitioners to move beyond evaluating LLMs in isolation, instead scrutinizing how biases can arise from or be amplified by the documents involved in RAG. Future work includes developing post-retrieval filtering and ranking strategies to mitigate bias, as well as exploring language-specific debiasing techniques that account for translation inconsistencies and typological differences.
@@
While this paper sheds light on how RAG affects social biases in LLMs, several important limitations warrant discussion. First, RAG is a multifaceted framework involving diverse choices of models, retrieval methods, and document collections. Although we explored a variety of LLMs, retrieval methods and datasets, our study did not encompass all possible combinations of these components, particularly those using domain-specific data or less common retrieval techniques due to the page limit. Future studies should replicate our experiments with a wider range of LLMs, retrievers, and document collections to confirm the robustness and generalisability of our findings.
Second, our analysis targeted three languages (i.e. English, Japanese, and Chinese) and four social bias types (i.e. gender, race, age, and religion). Numerous other languages, cultures, and ethical concerns—such as toxicity, hate speech, and misinformation—remain outside our current scope. Evaluating RAG systems for these additional dimensions is a critical step for achieving broader safety and fairness.
Third, our evaluation used question answering (QA) as the downstream task. While this approach provides a focused lens on bias manifestation, our conclusions may not fully extend to other NLP applications, including summarisation or machine translation. Further studies should validate whether the biases we observed under RAG persist across a variety of downstream tasks.
Lastly, although numerous techniques exist for debiasing LLMs (Li et al., 2024b; Lin et al., 2024; Li et al., 2024a), this paper did not systematically investigate how RAG interacts with those debiasing strategies. Exploring that interaction remains an open question and we encourage future work to assess whether debiasing methods can effectively mitigate biases arising from RAG.
@@
The document collections used for RAG are derived from publicly available sources as detailed in Table 1, where each dataset’s original authors labeled documents by bias type. Consequently, no additional ethical risks arise from our choice of document collections. Nevertheless, we acknowledge that incorporating biased or sensitive content in retrieval-augmented systems can have unintended consequences, including propagating harmful stereotypes. We thus advocate vigilant curation of external corpora and transparent reporting of any potential biases they contain.

### v2

@@ Abstract
Llm-jp-13B 17.96 / 6.55 23.02 / 15.67* 3.08* / 2.58 6.35* / -0.79
Qwen-3B 28.27 / 8.13 39.83* / 8.13 24.70 / -1.59 11.81* / -6.15*
Qwen-3B-Inst. 17.41 / 0.20 23.86 / 4.07 15.18 / -5.75 6.35* / -8.93*
Qwen-7B 18.85 / -1.39 27.88* / 0.00 17.91 / -3.97 10.02* / -8.63
Qwen-7B-Inst. 10.02 / -3.67 24.01* / 0.50 15.43 / -2.08 10.17 / -10.12
Qwen-14B 3.77 / -7.34 13.99* / -2.68 0.55 / -4.66 -4.51 / -8.93
Qwen-14B-Inst. -2.38 / -2.38 4.61 / 2.68 -3.08 / -5.95 -8.43 / -12.70*
Qwen-72B 9.58 / -3.89 22.12* / -3.08 11.26 / -3.08 9.33 / -3.97
Qwen-72B-Inst. -0.45 / 1.19 12.21* / -0.50 5.80 / -4.76 0.79 / -5.56*
o4-mini 2.12/ -4.56 9.67* / -3.77 7.24 / -4.86 4.81/ -4.26
@@
Table 2: Diff-Bias for the ambiguous and disambiguated gender contexts for different generator LLMs. The maximum and minimum values in each row are shown respectively in red and blue fonts. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
§ 4.3 Effect of Generators
While our main findings hold true on average across models (Figure 3), the choice of generator LLM introduces significant difference. We investigate three key aspects: the impact of instruction tuning, model scale, and specialised architectures like reasoning models. For this focused analysis, we present results on gender bias from the English BBQ dataset (see Table 2).
@@
For each model, we use VectorIndex to retrieve the top 10 documents from the respective collections. Table 2 shows LLMs trained on multilingual pre-train data in the top block, while models that are trained on increased proportions of Japanese and Chinese language pre-train data are shown respectively in the middle and bottom blocks.
Overall, every model exhibits increased gender bias when retrieving from the stereo-set, and decreased bias when retrieving from the anti-set. These findings corroborate the trend noted in Figure 3. We also find instruction tuning consistently reduces the baseline bias (w/o RAG). For instance, Qwen-14B-Inst. shows a significantly lower baseline bias than Qwen-14B. Similarly, larger instruct-based models generally exhibit lower intrinsic bias, with Qwen-72B-Inst. being the least biased in the w/o RAG setting. Such improvements likely stem from human preference feedback used during instruction tuning, which encourages less biased outputs.
However, this improved safety alignment is fragile and easily compromised by RAG. Table 2 shows that even the largest and best-aligned models still suffer a substantial increase in bias. This underscores that simply using a larger or better-aligned model is not a sufficient defence against bias injection from external documents.
To investigate whether models designed for complex reasoning are more robust, we evaluated o4-mini. While it demonstrates high accuracy across all RAG settings (95%, see Table 13), its Diff-Bias score still increases by over 7 points when exposed to the stereo-set. This reinforces our claim that retrieval content, not just model architecture, increases bias amplification.
@@
Table 3: Comparison of ambiguous and disambiguated Diff-Bias when using different retrieval methods to retrieving documents from the full-set. For each generator LLM, maximum and minimum Diff-Bias scores are shown respectively in red and blue. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
§ 4.4 Effect of Retrievers
Choice of Retrieval Method:
We compare the three retrieval methods—VectorIndex, BM25, and Contriever—by by retrieving 10 documents from the full-set on the English BBQ gender dataset. As shown in Table 3, all three RAG retrievers consistently amplify bias compared to the w/o RAG baseline, confirming that the act of retrieval itself is a significant source of social bias.
Furthermore, the Diff-Bias does not directly correlate with the proportion of stereotypical documents retrieved. For example, while the sparse retriever BM25 fetches the lowest percentage of stereotypical documents (46.04%), it induces the highest Diff-Bias across multiple generators (e.g., GPT-3.5, Qwen-14B). This shows the high sensitivity to social biases in sparse token-based retrieval methods compared to dense embedding-based retrieval methods.
Number of Retrieved Documents:
We then analyse the impact of varying the number of retrieved documents ($k$), presenting the average results across three LLMs with different model types, scales and families in Figure 4. The results show a trade-off and non-linear relationship. In ambiguous contexts (Figure 4, left), the Diff-Bias rises sharply with just a few documents (k=3), confirming that even minimal retrieval can significantly amplify bias.
However, as more documents are retrieved ($k>5$), the Diff-Bias begins to steady or even decrease, particularly for the full-set and anti-set. This suggests a potential trade-off between relevance and Diff-Bias (see Appendix E). This finding challenges the simple assumption that more stereo retrieved documents always leads to higher Diff-Bias.
3510203005510101515Diff-BiasAmbiguous Contexts
3510203010-108-86-64-42-20Disambiguated Contexts
Figure 4: Effect of different numbers of retrieved documents on Diff-Bias (averaged over Llama3-8B-Inst., Qwen-7B-Inst., Qwen-14B). X-axis shows the number of retrieved documents. Gray=w/o RAG, red=stereo-set, teal=full-set, blue=anti-set. The detailed figures for each model are shown in Figure 9 and Figure 10
Bias Type w/o RAG stereo-set full-set anti-set
CBBQ
Gender 8.96 / 5.41 31.76* / 16.81* 11.41 / 7.59 -8.33* / -1.21
Age 15.68 / -1.43 20.81* / 4.55 14.29 / -0.82 -0.28* / -3.50
Race 1.79 / 4.55 8.10* / 1.17 5.55 / -0.52 5.12 / -0.63
Religion 1.08 / 1.62 9.85* / -3.12 4.18 / -7.25 -1.70 / -7.25
JBBQ
Gender 1.50 / -9.83 11.22* / -7.32 6.67 / -7.42 4.28 / -9.12
Age 24.64 / -1.84 23.79 / 0.33 17.35 / -1.62 5.94* / -3.22
Table 4: Average Diff-Bias scores for CBBQ and JBBQ. Scores show the mean ambiguous / disambiguated Diff-Bias, averaged across GPT-3.5, Qwen-7B-Inst., Qwen-14B, Qwen-14B-Inst., and Qwen-72B-Inst. For each bias type, maximum and minimum Diff-Bias scores are shown respectively in red and blue. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
§ 4.5 Multilingual Bias Evaluation
To investigate our findings’ generalizability, we evaluate on Chinese (CBBQ (Huang and Xiong, 2024)) and Japanese (JBBQ (Yanaka et al., 2024)) datasets. As these languages lack dedicated stereotypical and anti-stereotypical corpora, we created document collections by machine-translating our English corpus. We also do human evaluation on the quality of translation in Appendix B.  [^2] Race and Religion categories are unavailable in JBBQ.
As shown in Table 4, across both languages, our main finding holds: RAG with a stereo-set consistently amplifies bias in ambiguous contexts, while the anti-set effectively reduces it compared to the stereo-set. However, we can also observe several dataset-specific differences as described next.
In CBBQ, while the amplification trend holds for ambiguous contexts, the disambiguated Religion category exhibits a counter-intuitive negative Diff-Bias. A model-by-model breakdown (Table 18) reveals that instruction-tuned Qwen2.5 models possess a strong intrinsic positive bias (e.g., +24.17), whereas the base Qwen2.5-14B has a strong intrinsic negative bias (-16.67). Furthermore, we find a distributional mismatch: 81.5% of CBBQ’s disambiguated religious questions feature Christians as the bias target group, while the stereo-set contains only approx. 9% Christian-related stereotypes. This distributional mismatch reduces the self-confidence in their initial biased answers. This finding highlights that RAG’s performance on mitigating or amplifying bias is highly sensitive to the distributional alignment between the query context and the knowledge corpus.
The bias amplification trend for the ambiguous contexts is consistent for Japanese as seen from the results on JBBQ. Moreover, the analysis presented in Appendix B indicates that machine translation sometimes fail to preserve certain nuances of the original stereotypes, and Japanese-specific issues such as zero-pronoun resolution Isozaki and Hirao (2003) (i.e. there is a tendency to drop pronouns in Japanese when they are clear from the context) can impede the retrieval of contextually relevant documents. These findings show the importance of creating high-quality multilingual bias corpora.
Model w/o RAG Stereo-set Full-set Anti-set
self-RAG 6.95 / 2.29 17.61* / 3.47 15.68 / 2.68 16.47 / 2.48
Vanilla 10.02 / -3.67 24.01* / 0.50 15.43 / -2.08 10.17 / -10.12
Table 5: Comparison of self-RAG against the vanilla base model Qwen-7B-Inst.. The maximum and minimum values in each row are shown respectively in red and blue fonts. 95 % CIs that do not overlap with the corresponding w/o RAG setting are indicated by *.
Model w/o RAG w/ RAG
Default ICL DDP Default ICL Summarizer DDP
GPT-3.5 5.16 / -9.33 15.43 / 4.37 -1.59 / -3.57 14.53 / 7.14 25.20 / 9.82 6.38 / -8.73 0.0 / -0.60
Qwen-7B-Inst. 10.02 / -3.67 9.38 / 7.44 0.69 / 3.57 24.01 / 0.50 21.78 / 6.85 18.30 / 1.19 23.0 / -1.39
Qwen-14B 3.77 / -7.34 3.32 / -3.76 3.67 / 2.68 13.99 /-2.68 13.63 / 7.04 9.52 / -3.89 20.44/ -0.40
Qwen-14B-Inst. -2.38 /-2.38 -1.14 / -5.46 3.27 / 1.19 4.61 / 2.68 2.43 / 1.88 -3.42 / 3.57 16.76 / 5.16
Table 6: Diff-Bias scores for the ambiguous and disambiguated contexts (values separated by ‘/’) under different debiasing strategies. In each group (“w/o RAG” and “w/ RAG”), for ambiguous and disambiguated values separately, the diff-bias with the lowest absolute value is highlighted in bold.
§ 4.6 Analysis of Advanced RAG
While advanced RAG frameworks like Self-RAG (Asai et al., 2024) and RankRAG (Yu et al., 2024) often outperform standard RAG systems, their integrated and complex structure presents challenges for social bias analysis. It remains unclear whether they can mitigate biases in RAG.
The original Self-RAG model uses Llama-2 as the base model, which frequently refuses to answer social bias-related questions. This is problematic when evaluating social biases in RAG. Therefore, following the original training procedure and training data, we trained Self-RAG version using Qwen-7B-Inst. We then compare this Self-RAG version against the vanilla Qwen-7B-Inst. model.
Table 5 shows that in the w/o RAG setting, Self-RAG demonstrates a lower Diff-Bias score compared against the vanilla model, suggesting that its self-correction and adaptive retrieval mechanisms are effective at reducing the social biases in the baseline generator. However, when retrieving documents from the stereo-set, Self-RAG’s Diff-Bias score increases to 17.61, confirming that Self-RAG is also affected by the bias amplification in RAG. Furthermore, unlike the vanilla model, Self-RAG shows high Diff-Bias scores not only on the full-set but also on the anti-set, where the vanilla model’s bias is significantly lower.
§ 4.7 Mitigating the Social Biases in RAG
Our analysis has revealed that both standard and even advanced RAG architectures like Self-RAG are highly susceptible to bias amplification by the retrieved documents. This underscores the critical need for effective mitigation strategies that can function within a RAG pipeline. In this section, we evaluate two categories of debiasing methods: (1) in-context prompting that guide model behaviour at inference time, and (2) alignment fine-tuning via DPO to reshape the model’s underlying preferences.
4.7.1 Prompting Strategies
Default: The model receives the original question without any additional prompts.
In-context Learning (ICL): Oba et al. (2024) used ICL to suppress biases in LLMs. We provide an LLM with eight curated few-shot examples from the English BBQ dataset to guide it towards unbiased generation (Oba et al., 2024). The examples cover ambiguous and disambiguated scenarios and are prepended to the user’s query (see Figure 5).
Summariser: To directly address stereotypes in the retrieved content, we use a summariser agent to distil the documents before they are passed to the generator. We use GPT-3.5-turbo with a prompt instructing it to produce a concise, neutral summary free of subjective or stereotypical word usages (Hu et al., 2024).
Dual Directional Prompting (DDP): We adapt the two-stage prompting strategy from Li et al. (2024a), where social attributes are first neutralised with placeholders for an initial prediction, which is then used to guide the final answer with an explicit instruction to ignore stereotypes.
As shown in Table 6, in the w/o RAG setting, both ICL and especially DDP are effective at mitigating the models’ intrinsic biases. For instance, DDP reduces the Diff-Bias of Qwen-7B-Inst. on ambiguous questions from 10.02 to nearly zero (0.69). However, in the w/o RAG setting, the Diff-Bias scores for ICL and DDP show little improvement over the default RAG baseline. This reveals a limitation of prompt-level strategies – the strong context injected by retrieved stereotypical documents can overwhelm prompt-based instructions or few-shot examples. In contrast, the summariser consistently reduces bias across all models in the RAG settings. Unlike other methods, the summariser directly tackles the biases from the retrieved documents before they influence the generator.
4.7.2 DPO Training
DPO (Rafailov et al., 2023) is an alignment algorithm that refines a model on preference data, using a “chosen” (preferred) and “rejected” response for each input. We apply DPO to investigate if reshaping a model’s preferences offers a more robust debiasing solution for RAG. Specifically, we fine-tune Qwen2.5‑7B‑Inst. on the GenderAlign dataset (Zhang et al., 2024), which contains 8000 dialogues paired with preferred (less biased) and rejected (more biased) responses.
Table 7 shows that DPO training successfully reduces bias in all conditions. The DPO-trained model exhibits a lower intrinsic bias in the w/o RAG setting. More importantly, it consistently achieves lower Diff-Bias scores across all retrieval sets (stereo, full, and anti) for ambiguous questions. The improvement on the disambiguated questions is smaller, which to be expected as the model is correctly relying on the provided context. This shows that DPO can be considered as a potential debiasing strategy for RAG systems by realigning the model’s internal behaviour.
Model w/o RAG stereo-set full-set anti-set
Vanilla 10.02 / -3.67 24.01 / 0.50 15.43 / -2.08 10.17 / -10.12
DPO 6.25 / -3.47 21.68 / 0.69 12.39 / -3.37 7.64 / -10.62
Table 7: Diff‐Bias scores for ambiguous and disambiguated contexts comparing the raw vs DPO-trained models on different RAG settings. Diff-bias with the lowest absolute value is shown in bold.
@@ 4.7.2 DPO Training
We conducted a comprehensive study on how RAG influences social biases LLM on three languages and discovered that RAG amplifies social biases in LLMs when stereotypical documents collection are used for retrieval. We evaluate several potential mitigating strategies for RAG and show the effectiveness of DPO. We urge practitioners to move beyond evaluating LLMs in isolation, and consider a wholistic evaluation within RAG.
@@
While this paper sheds light on how RAG affects social biases in LLMs, several important limitations warrant discussion. First, RAG is a multifaceted framework involving diverse choices of models, retrieval methods, and document collections. Although we explored a variety of LLMs, retrieval methods and datasets, our study did not encompass all possible combinations of these components, particularly those using domain-specific data or less common retrieval techniques due to the limited scope in a conference publication. Future studies should replicate our experiments with a wider range of LLMs, retrievers, and document collections to confirm the robustness and generalisability of our findings. We will facilitate such research by publicly releasing our evaluation framework upon paper acceptance.
Second, our analysis targeted three languages (i.e. English, Japanese, and Chinese) and four social bias types (i.e. gender, race, age, and religion). We selected those languages because of the availability of social bias evaluation datasets and the availability of native speakers of those languages who could evaluate the translation quality of the document collections. Furthermore, we selected those four bias types due to the availability of documents in the public domain that can be used in our retrieval experiments. The numbers of documents covering each social bias type is not equal as can be seen from Table 1. For example, most datasets cover gender-related social bias types well, whereas only four datasets include age-related social biases. Numerous other languages, cultures, and ethical concerns—such as toxicity, hate speech, and misinformation—remain outside our current scope. Evaluating RAG systems for these additional dimensions is a critical step for achieving broader safety and fairness guarantees.
Third, our evaluation used question answering (QA) as the main downstream task. We used three benchmarks for social bias evaluation that are publicly available and created by native speakers familiar with the respective cultures. These multilingual bias evaluation benchmarks are created following the English BBQ framework, but culturally adapted and independently validated by their respective authors. While QA provides a focused lens on bias manifestation, our conclusions may not fully extend to other NLP applications, such as summarisation or machine translation. Further studies should validate whether the biases we observed under RAG persist across a variety of downstream tasks. Importantly, no social bias evaluation benchmark is perfect – given that they are annotated by a small set of annotators, reflecting their own stereotypical viewpoints. However, BBQ, JBBQ and CBBQ benchmarks are widely used in prior work evaluating social biases, making them ideal candidates for our RAG social bias evaluations.
Numerous techniques LLMs (Li et al., 2024c; Lin et al., 2024; Li et al., 2024b) have been proposed for debiasing LLMs. In § 4.7, we considered prompt- and DPO-alignment-based methods for mitigating gender biases in the generator LLMs in a RAG system according to the availability of human-created preference dataset. Our experiments showed that, although those methods can indeed reduce the social biases in the generator LLMs, they increase again when stereotypical documents are used for RAG. Evaluating the effectiveness of all debiasing methods proposed in the prior work for LLMs is beyond the scope of this paper, but remains an important task before deciding which debiasing method should be used with RAG.
@@
The document collections used for RAG are derived from publicly available sources as explained in Table 1, where each dataset’s original authors have annotated the documents for bias types. Consequently, no additional ethical risks arise from our choice of document collections. Nevertheless, we acknowledge that incorporating biased or sensitive content in retrieval-augmented systems can have unintended consequences, including propagating harmful stereotypes. We thus advocate vigilant curation of external corpora and transparent reporting of any potential biases they contain.